# Second- and third-order properties of multidimensional Langevin equations

### v11

@@ 3.2. Quantitative significance of the deterministic contribution
The next question we wish to consider is that of quantitative significance. How do we judge whether or not the deterministic contribution $\boldsymbol{\alpha}^{\mathsf{T}}\mathbf{x}$ to $y$ is significant? And how do we compare whether a fitted model is a good match to experimental data, at least to linear order? This question has been addressed in the case of the usual stationary variables: compare the values of the experimentally measured covariance functions to the theoretical model prediction, and if the difference is small compared to $\mathbf{C}$, it is declared a good fit. However, we cannot proceed in an analogous way because the measured quantities are differences $y(\tau)-y(0)$ whose expectations scale as $\mathcal{O}(\tau^{1})$ whereas root-mean-square values $\langle(y(\tau)-y(0))^{2}\rangle^{1/2}$ scale as $\mathcal{O}(\tau^{1/2})$. Hence, they are not comparable.
@@
where $\lambda>0$. The dimensionless combination involving $\alpha$ is $(\alpha/\lambda)\sqrt{D^{xx}/D^{yy}}$. We will explicitly arrive at this quantity by studying the function $\langle(y(\tau)-y(0))^{2}\rangle$. We introduce according to Eq. (65) the variable $z$, with:
@@
From the covariance function above, we obtain:
@@ 3.3. Comparison of experimental data with theoretical predictions
To compare experimentally measured covariance functions to theoretical predictions, we can again normalize by the diffusion matrix. We may consider coordinates in which $\mathbf{D}=\mathds{1}$ and compare the experimental and theoretical values of $\langle\dot{\mathbf{x}}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$. Or, for a coordinate-independent description, similarly to the case of stationary variables, we may decompose the deviation of experimental to theoretical $\langle\dot{\mathbf{x}}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ into symmetric and antisymmetric parts, multiply on the right by $\mathbf{D}^{-1}$ and inspect the eigenvalues and compare them to unity. In terms of “ensemble variance”, this amounts to:
@@
where $\Delta_{\textrm{exp\textendash theo}}$ denotes the deviation of experimental to theoretical values. As mentioned for the angular momentum, if the time-step is not small, then on the r.h.s. of the above, it may be suitable to use discrete-time estimators of the diffusion. In case the eigenvalues of $\mathbf{H}$ exceed unity, the eigenvalues need to be compared to these instead, not just for the antisymmetric part but also for the symmetric part. We may illustrate this by way of two-dimensional linear Gaussian dynamics with complex conjugate eigenvalues and $D^{xx}=D^{yy}$, $D^{xy}=0$, in which case:
@@
where we re-indexed $m=n+k$. Thus, although the estimate Eq. (86) for the velocity autocovariance function is self-averaging for any given finite time lag, its sum over all time lags is not [^10] In the case where $\langle\Delta x\rangle$ is treated as known and can be used in place of $\overline{\Delta x}$ in Eq. (88), the resulting estimate of long-time diffusivity is $(N/2\Delta t)(\overline{\Delta x}-\langle\Delta x\rangle)^{2}$, which has the correct expectation value, but is not self-averaging.. If we denote the limits of summation by ±$\pm M$, then the limits $M\to\infty$, $N\to\infty$ do not commute.
@@ 4. Third-order properties
vanishes in the case where $\mathbf{x}$ has continuous paths [^11] The $x^{i}$, $x^{j}$, and $x^{k}$ need not be state coordinates; they can be any functions.. In such a situation, the probability density evolution satisfies a Fokker–Planck equation [3], leading to vanishing of the “cubic variation”. This gives a relation between the angular momenta, as:
@@
The angular momenta are related to probability currents. We start with the Fokker–Planck equation:
@@ 5. Inhomogeneous diffusion
In this section, we shall restrict ourselves to the case where $\langle\dot{\mathbf{x}}\mid\mathbf{x}\rangle$ is linear. In addition to being easier to deal with, in the case of weak nonlinearity, coordinates can be chosen via Koopman eigenfunctions [18, 19] to transform the system into this form, as will be done in the next section. Additional terms in inhomogeneous diffusion are assumed to be small.
@@
The second task is to compute covariance functions. In the current model, $\langle\mathbf{x}(\tau)\mid\mathbf{x}(0)\rangle=e^{\mathbf{A}\tau}\mathbf{x}(0)$, so by the law of iterated expectations:
@@
The question is, if this quantity is 0 for all $i,j,k$, are the covariance functions time-reversible? We see that we need the lower-order time-reversibility to hold, i.e., $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$. Calling $\widetilde{\mathbf{A}}(\tau)$ the long quantity in square brackets in Eq. (108) involving $\mathbf{A}$, we have the covariance function:
@@
Now in $\langle x^{i}(0)x^{j}(0)x^{k}(\tau)\rangle-\langle x^{i}(\tau)x^{j}(\tau)x^{k}(0)\rangle$ we can pull out a factor of $e^{\mathbf{A}\tau}\otimes e^{\mathbf{A}\tau}\otimes\mathds{1}-\mathds{1}\otimes\mathds{1}\otimes e^{\mathbf{A}\tau}$ and thus the condition for this quantity to be 0 for all $\tau$ and all $i,j,k$ reduces to that of $L(x^{i}x^{j},x^{k})=0$.
So far, we have understood how the quantities $b$ give rise to the quantities $\langle x^{i}x^{j}x^{k}\rangle$ and $L(x^{i}x^{j},x^{k})$. The remaining degree of freedom, $\langle x^{i}x^{j}\dot{x}^{k}\rangle$ is accounted for by nonlinearities, considered in the next section.
@@
Now, we address the case of integrated variables. Following the previous section, we assume such a variable $y$ obeys the law $\langle\dot{y}\mid\mathbf{x}\rangle=0$. We now use the model:
@@
If $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$, then calling $\widetilde{\mathbf{A}}(\tau)$ the long quantity in the square brackets involving $\mathbf{A}$:
@@ 6.2. One-dimensional case
Although the dynamics of $\mathbf{x}$ may have arbitrary non-linearities, the Koopman operator is a linear operator. The significance of the eigenfunctions of the Koopman operator (hereafter called “Koopman eigenfunctions”) is that their dynamics follows a linear law in expectation. (Techniques and results for estimation of the Koopman operator are presented in [20, 21].) The Koopman operator $\mathcal{K}$ has the same eigenvalues as the Perron–Frobenius operator $\mathcal{P}$ appearing in the Fokker–Planck equation ($\partial p(\mathbf{x},t)/\partial t=\mathcal{P}p(\mathbf{x},t)$, where $\mathcal{P}$ operates on $\mathbf{x}$), because $\mathcal{K}$ is the adjoint of $\mathcal{P}$ [19] with respect to the inner product defined by:
@@
Addressing the other covariance function, in one dimension this should be identical, but we calculate it again here. The eigenfunction $f_{2}(x(\tau))$ involves ${He}_{3}(x(\tau))$, but when multiplying by $x(0)$ and taking the expectation, it vanishes to zeroth order. Thus Isserlis’s theorem holds again to zeroth order. The result is:
@@ 6.3. Multidimensional case
For the covariance functions, we solve a differential equation, replacing fourth-order moments using Isserlis’s theorem, with $\mathcal{O}((a,b)^{2})$ corrections. Without going through the details of the calculations, we use again the quantity we called $\widetilde{\mathbf{A}}(\tau)$ (the long quantity in square brackets in Eq. (108)), and the result is:
@@
At this point, we note that by a similar calculation, if there are cubic terms in the drift or quadratic terms in the diffusivity, then the corrections to the second- and third-order covariance functions are quadratic in the coefficients deviating from a linear Gaussian model (written in terms of multidimensional Hermite polynomials). Beyond this, the calculations may not work as expected (see Appendix B).
@@ 6.7. Finite time interval inhomogeneous diffusion
When the variables are of even parity (i.e., invariant under time reversal), the coefficients describing inhomogeneous diffusion are time-symmetric and can be estimated by way of the time-symmetric quantities:
@@ 7.1. Linear Gaussian systems
(216) $\dot{\mathbf{x}}^{+}$ $=\mathbf{A}^{+}_{+}\mathbf{x}^{+}+\mathbf{A}^{+}_{-}\mathbf{x}^{-}+\boldsymbol{\xi}^{+},$
(217) $\dot{\mathbf{x}}^{-}$ $=\mathbf{A}^{-}_{+}\mathbf{x}^{+}+\mathbf{A}^{-}_{-}\mathbf{x}^{-}+\boldsymbol{\xi}^{-},$
(218) $\langle\boldsymbol{\xi}^{\pm}(t)\rangle=\mathbf{0},\quad\langle\boldsymbol{\xi}^{\pm}(t){\boldsymbol{\xi}^{\pm}(t^{\prime})}^{\mathsf{T}}\rangle=2\mathbf{D}^{\pm\pm}\delta(t-t^{\prime}),\quad\langle\boldsymbol{\xi}^{+}(t){\boldsymbol{\xi}^{-}(t^{\prime})}^{\mathsf{T}}\rangle=2\mathbf{D}^{+-}\delta(t-t^{\prime}),$
where $\boldsymbol{\xi}^{\pm}$ are Gaussian random noise and all coefficients are constant. To have time-reversal symmetry, we must have $\mathbf{D}^{+-}=\mathbf{0}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle=\mathbf{0}$. These imply:
@@
Next, we address the time-reversibility of the second moments. Similarly as before, this depends only on the linear dynamics and not on the homogeneity of the diffusion matrix.
@@
(230) $\langle\xi^{y}(t)\rangle=0,\quad\langle\boldsymbol{\xi}^{\pm}(t)\xi^{y}(t^{\prime})\rangle=2\mathbf{D}^{y\pm}\delta(t-t^{\prime}),\quad\langle\xi^{y}(t)\xi^{y}(t^{\prime})\rangle=2D^{yy}\delta(t-t^{\prime}),$
where $\xi^{y}$ is Gaussian random noise and all coefficients are constant. For time-reversibility, we need $\mathbf{D}^{y-}=\mathbf{0}$. We write $\Delta y(\tau):=y(\tau)-y(0)$, and:
@@
where we labeled $\mathbf{x}^{\pm}\Delta y$ as $y^{(1)\pm}$. Again, we write $\widehat{\mathbf{A}}$ for the matrix on the r.h.s. and the vector [^15] A collection of quantities, not a true physical vector obeying the transformation laws under a change of basis. appearing in the above differential equation is equal to $e^{\widehat{\mathbf{A}}\tau}$ multiplied by its value when $\tau=0$. The elements are:
@@
which can be easily shown by induction provided that the $\mathbf{x}$ variables satisfy detailed balance. The case for an odd integrated variable is analogous.
@@ 7.2. Third-order quantities
Here we discuss the conditions for detailed balance in the Fokker–Planck equation Eq. (93) in the presence of odd variables. Each variable $x^{i}$ has a parity $\epsilon^{i}=\pm 1$, where under time-reversal $x^{i}\to\epsilon^{i}x^{i}$ ($i$ not summed over), and the product $\boldsymbol{\epsilon}\mathbf{x}$ is defined as $(\boldsymbol{\epsilon}\mathbf{x})^{i}=\epsilon^{i}x^{i}$ ($i$ not summed over), following [3]. In terms of the stationary distribution $p(\mathbf{x})$, the conditions for detailed balance are [3]:
(238) $\epsilon^{i}\mathcal{A}^{i}(\boldsymbol{\epsilon}\mathbf{x})p(\mathbf{x})$ $=-\mathcal{A}^{i}(\mathbf{x})p(\mathbf{x})+\partial_{j}[\mathcal{B}^{ij}(\mathbf{x})p(\mathbf{x})]$
(239) $\epsilon^{i}\epsilon^{j}\mathcal{B}^{ij}(\boldsymbol{\epsilon}\mathbf{x})$ $=\mathcal{B}^{ij}(\mathbf{x}).$
where $i$ or $j$ appearing twice because of $\boldsymbol{\epsilon}$ is not summed over. Now, as before we multiply Eq. (238) by a test function $f(\mathbf{x})$ and integrate. We consider test functions that are either even or odd under time reversal, i.e., $f(\boldsymbol{\epsilon}\mathbf{x})=\pm f(\mathbf{x})$. We also have a necessary condition for detailed balance:
(240) $p(\boldsymbol{\epsilon}\mathbf{x})=p(\mathbf{x}),$
which means that we may write $\langle g(\boldsymbol{\epsilon}\mathbf{x})\rangle=\langle g(\mathbf{x})\rangle$ for any function $g(\cdot)$. If $f(\cdot)$ and $x^{i}$ have the same parity, then we arrive back to the same condition as the case considered earlier, i.e., where there are no odd variables. On the other hand, if $f(\cdot)$ and $x^{i}$ have the opposite parity, then upon multiplying by $f(\mathbf{x})$ and integrating, the l.h.s. of Eq. (238) cancels the first term on the r.h.s. of Eq. (238), and the remaining term reads:
@@
which is guaranteed by substituting $\mathbf{x}\to\boldsymbol{\epsilon}\mathbf{x}$ (allowed because of Eq. (240)) and applying Eq. (239). Thus Eq. (238) gives no further information besides that contained in Eq. (239) and Eq. (240) in the case where odd quantities are considered.

### v12

@@ 3.2. Quantitative significance of the deterministic contribution
The next question we wish to consider is that of quantitative significance. How do we judge whether or not the deterministic contribution $\bm{\alpha}^{\mathsf{T}}\mathbf{x}$ to $y$ is significant? And how do we compare whether a fitted model is a good match to experimental data, at least to linear order? This question has been addressed in the case of the usual stationary variables: compare the values of the experimentally measured covariance functions to the theoretical model prediction, and if the difference is small compared to $\mathbf{C}$, it is declared a good fit. However, we cannot proceed in an analogous way because the measured quantities are differences $y(\tau)-y(0)$ whose expectations scale as $\mathcal{O}(\tau^{1})$ whereas root-mean-square values $\langle(y(\tau)-y(0))^{2}\rangle^{1/2}$ scale as $\mathcal{O}(\tau^{1/2})$. Hence, they are not comparable.
@@
where $\lambda>0$. The dimensionless combination involving $\alpha$ is $(\alpha/\lambda)\sqrt{D^{xx}/D^{yy}}$. We will explicitly arrive at this quantity by studying the function $\langle(y(\tau)-y(0))^{2}\rangle$. We introduce according to Eq. (66) the variable $z$, with transformed diffusion coefficients $D^{xz}$ and $D^{zz}$:
@@
From the covariance function in the previous subsection, we obtain:
@@ 3.3. Comparison of experimental data with theoretical predictions
To compare experimentally measured covariance functions to theoretical predictions, we can again normalize by the diffusion matrix. We may consider coordinates in which $\mathbf{D}=\mathds{1}$ and compare the experimental and theoretical values of $\langle\dot{\mathbf{x}}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$. Or, for a coordinate-independent description, similarly to the case of stationary variables, we may decompose the deviation of experimental to theoretical $\langle\dot{\mathbf{x}}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ into symmetric and antisymmetric parts, multiply on the right by $\mathbf{D}^{-1}$ and inspect the eigenvalues and compare them to unity. In terms of “ensemble covariance”, this amounts to:
@@
where $\Delta_{\textrm{exp\textendash theo}}$ denotes the deviation of experimental to theoretical values. As mentioned for the angular momentum, if the time-step is not small, then on the r.h.s. of the above, it may be suitable to use discrete-time estimators of the diffusion. In case the eigenvalues of $\mathbf{H}$ exceed unity, the eigenvalues need to be compared to these instead, not just for the antisymmetric part but also for the symmetric part. We may illustrate this by way of two-dimensional linear Gaussian dynamics with complex conjugate eigenvalues $-\lambda\pm i\omega$ of $\mathbf{A}$ (Eq. (35)) and $D^{xx}=D^{yy}$, in which case:
@@
where we re-indexed $m=n+k$. Thus, although the estimate Eq. (87) for the velocity autocovariance function is self-averaging for any given finite time lag, its sum over all time lags is not [^10] In the case where $\langle\Delta x\rangle$ is treated as known and can be used in place of $\overline{\Delta x}$ in Eq. (89), the resulting estimate of long-time diffusivity is $(N/2\Delta t)(\overline{\Delta x}-\langle\Delta x\rangle)^{2}$, which has the correct expectation value, but is not self-averaging.. If we denote the limits of summation by ±$\pm M$, then the limits $M\to\infty$, $N\to\infty$ do not commute. The same considerations apply when integrating the covariance functions of stationary variables.
@@ 4. Third-order properties
vanishes in the case where $\mathbf{x}$ has continuous paths [^11] The $x^{i}$, $x^{j}$, and $x^{k}$ need not be state coordinates; they can be any functions.. In such a situation, the probability density evolution satisfies a Fokker–Planck equation [3], leading to vanishing of the “cubic variation” (Eq. (92)) [18]. This gives a relation between the angular momenta, as:
@@
The angular momenta are related to probability currents. We start with the Fokker–Planck equation for the probability density $p(\mathbf{x},t)$ corresponding to the Itô–Langevin equation Eq. (1):
@@ 5. Inhomogeneous diffusion
In this section, we shall restrict ourselves to the case where the drift $\bm{\mathcal{A}}(\mathbf{x})$ is a linear function $\mathbf{A}\mathbf{x}$. In addition to being easier to deal with, in the case of weak nonlinearity, coordinates can be chosen via Koopman eigenfunctions [19, 20] to transform the system into this form, as will be done in the next section. Additional terms in inhomogeneous diffusion are assumed to be small.
@@
The second task is to compute covariance functions. In the current model, $\langle\mathbf{x}(\tau)\mid\mathbf{x}(0)\rangle=e^{\mathbf{A}\tau}\mathbf{x}(0)$ as derived before, so by the law of iterated expectations:
@@
The question is, if this quantity is 0 for all $i,j,k$, are the covariance functions time-reversible? We will see that we need the lower-order time-reversibility to hold, i.e., $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$. Calling $\widetilde{\mathbf{A}}(\tau)$ the quantity in square brackets in Eq. (109), we have the covariance function:
@@
Now in $\langle x^{i}(0)x^{j}(0)x^{k}(\tau)\rangle-\langle x^{i}(\tau)x^{j}(\tau)x^{k}(0)\rangle$ we can pull out a factor of $e^{\mathbf{A}\tau}\otimes e^{\mathbf{A}\tau}\otimes\mathds{1}-\mathds{1}\otimes\mathds{1}\otimes e^{\mathbf{A}\tau}$ and thus the condition for this quantity to be 0 for all $\tau$ and all $i,j,k$ reduces to $L(x^{i}x^{j},x^{k})=0$.
So far, we have understood how the quantities $b$ give rise to the quantities $\langle x^{i}x^{j}x^{k}\rangle$ and $L(x^{i}x^{j},x^{k})$. The remaining degree of freedom, $\langle x^{i}x^{j}\dot{x}^{k}\rangle$ is accounted for by nonlinearities in the drift, considered in the next section.
@@
Now, we address the case of integrated variables. We again assume a linear drift function. Following the previous section, we can make a transformation so that $y$ obeys the law $\langle\dot{y}\mid\mathbf{x}\rangle=0$. We now use the model:
@@
If $\mathbf{A}\mathbf{C}=\mathbf{C}\mathbf{A}^{\mathsf{T}}$, then calling $\widetilde{\mathbf{A}}(\tau)$ the quantity in square brackets above:
@@ 6.2. One-dimensional case
Although the dynamics of $\mathbf{x}$ may have arbitrary nonlinearities, the Koopman operator is a linear operator. The significance of the eigenfunctions of the Koopman operator (hereafter called “Koopman eigenfunctions”) is that their dynamics follows a linear law in expectation. (Techniques and results for estimation of the Koopman operator are presented in [21, 22].) The Koopman operator $\mathcal{K}$ has the same eigenvalues as the Perron–Frobenius operator $\mathcal{P}$ appearing in the Fokker–Planck equation ($\partial p(\mathbf{x},t)/\partial t=\mathcal{P}p(\mathbf{x},t)$, where $\mathcal{P}$ operates on $\mathbf{x}$), because $\mathcal{K}$ is the adjoint of $\mathcal{P}$ [20] with respect to the inner product defined by:
@@
Addressing the other covariance function $\langle x(\tau)^{2}x(0)\rangle$, in one dimension this is necessarily identical because the probability current vanishes [18] implying time-reversal invariance [3], but we explicitly calculate it here for illustration. The eigenfunction $f_{2}(x(\tau))$ involves ${He}_{3}(x(\tau))$, but when multiplying by $x(0)$ and taking the expectation, it vanishes to zeroth order. Thus Isserlis’s theorem holds again to zeroth order. The result is:
@@ 6.3. Multidimensional case
For the covariance functions, we solve a differential equation, replacing fourth-order moments using Isserlis’s theorem, with $\mathcal{O}((a,b)^{2})$ corrections. Without going through the details of the calculations, we use again the quantity in square brackets in Eq. (109), which we call $\widetilde{\mathbf{A}}(\tau)$, and the result is:
@@
At this point, we note that by a similar calculation, if there are cubic terms in the drift or quadratic terms in the diffusivity, then the corrections to the second- and third-order covariance functions are quadratic in these coefficients (multiplying multidimensional Hermite polynomials). Beyond this, the calculations may not work as expected (see Appendix B).
@@ 6.7. Finite time interval inhomogeneous diffusion
The coefficients describing inhomogeneous diffusion can be estimated by way of the quantities:
@@ 7.1. Linear Gaussian systems
(217) $\dot{\mathbf{x}}^{+}$ $=\mathbf{A}^{+}_{+}\mathbf{x}^{+}+\mathbf{A}^{+}_{-}\mathbf{x}^{-}+\bm{\xi}^{+},$
(218) $\dot{\mathbf{x}}^{-}$ $=\mathbf{A}^{-}_{+}\mathbf{x}^{+}+\mathbf{A}^{-}_{-}\mathbf{x}^{-}+\bm{\xi}^{-},$
(219) $\quad\langle\bm{\xi}^{\pm}(t){\bm{\xi}^{\pm}(t^{\prime})}^{\mathsf{T}}\rangle=2\mathbf{D}^{\pm\pm}\delta(t-t^{\prime}),\quad\langle\bm{\xi}^{+}(t){\bm{\xi}^{-}(t^{\prime})}^{\mathsf{T}}\rangle=2\mathbf{D}^{+-}\delta(t-t^{\prime}),$
where $\bm{\xi}^{\pm}$ are zero-mean Gaussian white noise and all coefficients are constant. To have time-reversal symmetry, we must have $\mathbf{D}^{+-}=\mathbf{0}$ and $\langle\mathbf{x}^{+}{\mathbf{x}^{-}}^{\mathsf{T}}\rangle=\mathbf{0}$. These imply:
@@
Next, we address the time-reversibility of the second moments. Similarly as before, this depends only on the linearity of the drift and not on the homogeneity of diffusion. For time-reversibility, we must have:
@@
(231) $\langle\xi^{y}(t)\bm{\xi}^{\pm}(t^{\prime})\rangle=2\mathbf{D}^{y\pm}\delta(t-t^{\prime}),\quad\langle\xi^{y}(t)\xi^{y}(t^{\prime})\rangle=2D^{yy}\delta(t-t^{\prime}),$
where $\xi^{y}$ is zero-mean Gaussian white noise and all coefficients are constant. For time-reversibility, we need $\mathbf{D}^{y-}=\mathbf{0}$. We write $\Delta y(\tau):=y(\tau)-y(0)$, and:
@@
where we labeled $\mathbf{x}^{\pm}\Delta y$ as $y^{(1)\pm}$. Again, we write $\widehat{\mathbf{A}}$ for the matrix on the r.h.s. and the vector [^15] An ordered collection of quantities, not a true physical vector obeying the transformation laws under a change of basis. appearing in the above differential equation is equal to $e^{\widehat{\mathbf{A}}\tau}$ multiplied by its value when $\tau=0$. The elements are:
@@
which can be easily shown by induction provided that the $\mathbf{x}$ variables satisfy detailed balance. This establishes the time-reversibility of $\langle(y(\tau)-y(0))\mathbf{x}(0)\rangle$. The case for an odd integrated variable is analogous.
@@ 7.2. Third-order quantities
Here we discuss the conditions for detailed balance in the Fokker–Planck equation Eq. (94) in the presence of odd variables. Each variable $x^{i}$ has a parity $\epsilon^{i}=\pm 1$, where under time-reversal $x^{i}\to\epsilon^{i}x^{i}$ ($i$ not summed over), and the product $\bm{\epsilon}\mathbf{x}$ is defined as $(\bm{\epsilon}\mathbf{x})^{i}=\epsilon^{i}x^{i}$ ($i$ not summed over), following [3]. In terms of the stationary distribution $p(\mathbf{x})$, the conditions for detailed balance are [3]:
(239) $\epsilon^{i}\mathcal{A}^{i}(\bm{\epsilon}\mathbf{x})p(\mathbf{x})$ $=-\mathcal{A}^{i}(\mathbf{x})p(\mathbf{x})+\partial_{j}[\mathcal{B}^{ij}(\mathbf{x})p(\mathbf{x})]$
(240) $\epsilon^{i}\epsilon^{j}\mathcal{B}^{ij}(\bm{\epsilon}\mathbf{x})$ $=\mathcal{B}^{ij}(\mathbf{x}).$
where $i$ or $j$ appearing twice because of $\bm{\epsilon}$ is not summed over. Now, as before we multiply Eq. (239) by a test function $f(\mathbf{x})$ and integrate. We consider test functions that are either even or odd under time reversal, i.e., $f(\bm{\epsilon}\mathbf{x})=\pm f(\mathbf{x})$. We also have a necessary condition for detailed balance:
(241) $p(\bm{\epsilon}\mathbf{x})=p(\mathbf{x}),$
which means that we may write $\langle g(\bm{\epsilon}\mathbf{x})\rangle=\langle g(\mathbf{x})\rangle$ for any function $g(\cdot)$. If $f(\cdot)$ and $x^{i}$ have the same parity, then we arrive back to the same condition as the case where there are no odd variables. On the other hand, if $f(\cdot)$ and $x^{i}$ have the opposite parity, then upon multiplying by $f(\mathbf{x})$ and integrating, the l.h.s. of Eq. (239) cancels the first term on the r.h.s. of Eq. (239), and the remaining term reads:
@@
which is guaranteed by substituting $\mathbf{x}\to\bm{\epsilon}\mathbf{x}$ (allowed because of Eq. (241)) and applying Eq. (240). Thus Eq. (239) gives no further information besides that contained in Eq. (240) and Eq. (241) in the case where odd quantities are considered.